OpenStudy (anonymous):

sin(-150°) = -sin30° true or false?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Flase

OpenStudy (anonymous):

False*

OpenStudy (anonymous):

The identity \( \sin(-\theta) = -\sin(\theta) \)

OpenStudy (anonymous):

is that how you determine?

OpenStudy (anonymous):

Dude plug it into a calculator

OpenStudy (anonymous):

its not an identity its simply asking if they equal each other and even if it was an identity it still would be false

Vocaloid (vocaloid):

aren't they both equal to -1/2?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Here http://www.sosmath.com/trig/Trig5/trig5/trig5.html The answer is True

OpenStudy (anonymous):

And it is an identity

OpenStudy (anonymous):

\[\sin (-150)=.714 -\sin (30)=.9\]

OpenStudy (anonymous):

Its False

Vocaloid (vocaloid):

Nixy is correct.

OpenStudy (anonymous):

it is in radian

OpenStudy (anonymous):

I mean to degrees

Vocaloid (vocaloid):

The question is asking for degrees...

OpenStudy (anonymous):

We are working with degrees

OpenStudy (anonymous):

I am correct. It is an even odd identity and it is True.

Vocaloid (vocaloid):

 confirm

OpenStudy (anonymous):

Thank you everyone

OpenStudy (anonymous):

YW